# What is “quantum” about quantum gravity?

### v1

@@ Abstract
Assuming the validity of the equivalence principle in the quantum regime, we argue that one of the assumptions of the usual definition of quantum mechanics, namely separation between the “classical” detector and the “quantum” system, must be relaxed. We argue, therefore, that if both the equivalence principle and quantum mechanics continue to survive experimental tests, that this favors “epistemic” interpretations of quantum mechanics (where formalism is built around relations between observables) over “ontic ones” (assuming the reality of states and wavefunctions). In particular, we show that relational type interpretations can readily accomodate the equivalence principle via a minor modification of the assumptions used to justify the formalism. We qualitatively speculate what a full generally covariant quantum dynamics could look like, and comment on experimental investigations.
From the first attempts at quantizing gravity bronstein it became clear that there must be something fundamentally different about it w.r.t. the other quantum theories. The basic issue is the equivalence principle. As discussed in bronstein , there is tension between quantum canonical uncertainty relations,
@@
The main point is that the usual “mathematical apparatus of quantum mechanics”, describing states as vectors in Hilbert space and observables as operators of this space, essentially assumes that the detector just “projects out the operator corresponding to the observable”, and in the process “collapses the wavefunction”. Many believe these fundamental assumptions of quantum formalism to be “physical”, and this belief is tied to the belief that quantum states and wavefunctions are “objective physical objects” (see for instance wigner ). In such a case, Bronstein’s argument implies that, provided the equivalence principle is exactly valid, there is new physics beyond quantum mechanics when gravitational effects are present (perhaps of the form penrose ; curce ), whose mathematical formalization is wholly new, given the “tight” axiomatic structure of quantum theory witalg ; vicary .
However, others (generally ascribing to QBism, relationalism, the relative state and other “epistemic” formulations) interp think that states and wavefunctions are nothing more than convenient mathematical objects to parametrize knowledge accessible to us. After all, non-commuting operators are inevitable if one assumes that some measurements are inherently incompatible with others; Furthermore, if our theory is to accomodate that
@@
Hermitian operators and Hilbert spaces then become the mathematical tool with the required characteristics. In this philosophy, the issues Bronstein raised are ”technical” rather than fundamental ones. One would just have to find more general mathematics objects, with better-suited characteristics in the regime where detector backreaction is significant.
@@
Let us think what happens when the concepts above refer to coordinates, and one expects all quantum observables (i.e. correlators to all orders) to be generally covariant w.r.t. coordinate changes frames . It will mean that
@@
At first sight, this looks like a trivial generalization, but it is far from the case. In rovelli , the formalism is build around discreteness (ingrained in $N$ and the finite dimension of $U_{a,b}$). But coordinate transformations are continuus, necessitating a limit of $N\rightarrow\infty$ while $F$ become a function of the position operator $\hat{x}$ in quantum mechanics (one can define this via a lattice in configuration space). In quantum field theory, $F$ comes back to be a function of the label $x$ but $U$ becomes an operator valued function of the field operator $\hat{\phi}(x)$. As we know, the analytical properties of limits such as $N\rightarrow\infty$ can be different from any finite $N$ properties.
@@
associated with relativistic non-inertial frames. If the equivalence principle is exact, no contact with the regin beyond the horizon is possible. Hence,$F_{ij}$ is generally non-unitary and $U^{f},U^{f^{\prime}}$ are in different unitarity classes. Degrees of freedom beyond the horizon must be traced out by $F_{ij}$ for all questions, a la laflamme . Thus, transitivity as well as Eq. 4 fails and $F^{+}F\neq\delta$. Note that a time evolution constructed in a metric with a causal horizon should not be a unitary operator, but at most an operator of a Lindbladian type lind , transforming covariantly under diffeomorphism shifts that exchange causal foliations
@@
Thus, as described in witalg the observables accessible to a detector can not be described by a “wave-function” vector in Hilbert space if Eq. 5 holds. If however these observables form a type II algebra, as witalg argues for, they can be described by a “wave-function” as a parametrization of a non-pure density matrix, $\hat{\rho}_{ij}=\left|\Psi\right>_{i}\prescript{}{j}{\left<\Psi\right|}$ (but since there is no Hermitian Hamiltonian there are also no pure states commuting with it) or equivalently a partition function nishioka (where general convariance is imposed both on the action and the domain of functional integration, creating a ”holographic” constraint holo ). The question of what substitutes unitarity as a constraint on operators in the presence of general covariance is of course very much open. In megrav it is speculated that for general covariance unitarity $\left(\mathrm{Re}U+\mathrm{Im}U=\left(\mathrm{Re}U-\mathrm{Im}U\right)^{-1}\right)$ is relaxed to a general Kramers-Konig type tong fluctuation dissipation relation $\left(\begin{array}[]{cc}\mathrm{Re}\\ \mathrm{Im}\end{array}U=\begin{array}[]{c}F_{1}(\mathrm{Im}U)\\ F_{2}(\mathrm{Re}U)\end{array}\right)$ for some $F_{1,2}(...)$ determined locally by dynamics, where the ”Real” fluctuation term dominates for weak fields (and represents the equivalence between coordinate metric superpositons and detector superpositions) while the ”Imaginary” dissipation dominates for stronger fields and represents decoherence from horizons, a la wald .
@@
Thus, the problem of formulating a theory which is both quantum and obeys the equivalence principle exactly has been reformulated as a problem of finding a generally covariant representation for $F,U$ in Eq. 5. This is of course a very difficult problem ($F,U$ are infinite dimensional, and $F$ represent coordinate diffeomorphisms). So far it is fair to say no such theory is conclusively proven to exist. The only candidates of quantum gravity explicitly built around the fully quantum equivalence principle, Loop Quantum Gravity and causal dynamical triangulation lqg ; cdt identify spacetime coordinates (through line,area and volume operators) with observables (so there is no separation between $U$ and $F$) and impose general covariance on the state, something that is different from the relational discussion in this work, centered around observables rather than states. In realistic physical experiments we do not measure coordinates directly, but infer them via interactions (“rulers” and “clocks”), hence the distinction between $U$ and $F$. General govariance means not the covariance of a “state” but of observable correlations (matrices $\left|U_{ab}\right|^{2}$, whose basis are field correlators, under $F-$ type transformations), and general covariance for field correlators living on a quantum spacetime is not as yet proven for LQG or CDT lqg ; cdt . However, to tree level in certain limits quantum field theory already fulfills the requirement of general covariance matsas so the hope that this can be brought to higher orders in an EFT expansion, as suggested by megrav ; holo is not unfounded. This approach entails reconstructing the metric from QFT observables and seeing what sort of QFT action would lead to a general covariance in terms of this of metric. Such investigations are speculative but ongoing (eg padma ; kempf1 ; kempf2 ; tales ; nist ).
@@
Most importantly, the problem described in the above paragraph can be thought of as ”technical” rather than ”conceptual”, for a price which is itself however both technical and conceptual in nature. The technical price is the abandonment of the very convenient Hilbert space and Unitary/Hermitian operator technology. The conceptual price is a definite abandonment of notions that wavefunctions and states are physical objects. General covariance makes it inevitable that the only mathematical objects readily associated with a physical meaning are “sets of questions that can be asked of a system”, i.e. density operators in some observable’s basis. For anyone who does not associate an ontic truth to quantum states, this is a small price to pay, or equivalently a feature rather than a bug.
@@
Looking at this experiment from the point of view of quantum reference frames frames indeed confirms that the analysis in witness ; witness2 ; gyro ; gyro2 ; pend1 is done in the laboratory frame. Using the spin set-up in witness , Eq. 5’s $U$ will be in the basis of all possible $\uparrow_{1}\uparrow_{2}(x_{in})\right|\left|\uparrow_{1}\downarrow_{2}(x_{out})\right>$ where $\uparrow_{1,2}$ transforms as a spinor under the Lorentz group. Provided the nanoparticles being and end in the same place (something not explicitly required in witness ; witness2 but can be easily accomplished, as Fig. 1 shows) the phase change is the difference between two angles propagated through two proper time intervals. Thus in the classical limit binney given a unit vector $A^{\mu}$ the phase difference $\sim(\Delta A_{\mu})A^{\mu}$ where $\Delta A^{\mu}=R^{\mu}_{\alpha\beta\gamma}\Delta S^{\alpha\beta}A^{\gamma}$ ($A^{\mu}$ is the vector projection of the difference between two spinors, a vector), where $S^{\alpha\beta}$ is the area element parallel to the interferoemeter and $R^{\mu}_{\alpha\beta\gamma}$ the Riemann tensor of the gravitational field (sum of earth’s, apparatus and nanoparticles). If tidal forces were to be ignored (they can for the earth, but not the rest) this would of course be a 4-vector. We are looking, therefore, for the observable to transform covariantly as a 4-vector if general covariance works in the quantum regime.
@@
Thus the study of quantum gravity might be in it’s ”golden” phase: Deep conceptual issues in both gravity (the role of the equivalence principle) and quantum mechanics (fundamental limits to the “classicality” of the detector and the “quantumness of the system) might be directly tested by experiments of the type witness ; witness2 ; gyro ; gyro2 ; pend1 . If the answer estimated by witness will be seen, one can interpret this as a confirmation of the quantization of gravity but also an experimentally verified breakdown of general covariance of quantum observables (or rather of the vacuity of the concept of general covariance in the quantum regime), as advocated in dono . If the result of witness ; witness2 ; gyro ; gyro2 ; pend1 can not be described by canonical quantum mechanics but rather with some kind of mixed theory a la oppenheim ; megrav ; holo and at the same time stringent tests of the equivalence principle with both astrophysical observations nordveldt and tabletop measurements atoms yield no violation, one would have to think deeper at how to implement the equivalence principle into a “relational generally covariant Algebra of observables” woit ; witalg , which in general will not be based around unitary matrices. The advent of mesoscopic experimental study of quantum effects on gravity, therefore, promises to be important and fruitful.

### v2

@@ Abstract
Assuming the validity of the equivalence principle in the quantum regime, we argue that one of the assumptions of the usual definition of quantum mechanics, namely separation between the “classical” detector and the “quantum” system, must be relaxed. We argue, therefore, that if both the equivalence principle and quantum mechanics continue to survive experimental tests, that this favors ”informational” interpretations of quantum mechanics (where formalism is built around relations between observables, defined as information that can be accessed by an observer of the system. In particular ”collapse” is understood as a change of relative information as the detector interacts with the system ) over “ontic ones” (assuming the physical reality of states and wavefunctions, which are assumed to be more than informational objects. In particular collapse is understood as a physical process). In particular, we show that relational type interpretations can readily accomodate the equivalence principle via a minor modification of the assumptions used to justify the formalism. We qualitatively speculate what a full generally covariant quantum dynamics could look like, and comment on experimental investigations.
I Introduction
From the first attempts at quantizing gravity bronstein it became clear that there must be something fundamentally different about it w.r.t. the other quantum theories. The basic issue is the equivalence principle [^1] This term of course can mean several different things. For the considerations of most of this work the weak non-relativistic form, that the inertial and gravitational mass are equal, is sufficient. However Section II defines the equivalence principle as the general covariance of all observables (quantum in this case) which is a stronger form. As discussed in bronstein , there is tension between quantum canonical uncertainty relations,
@@ I Introduction
One could argue, as does a good part of the modern literature on the topic (the ”holographic approach”, where the gravitational degrees of freedom of a region of spacetime is always encoded on the boundary harlow , or the “S-matrix approach” where all gravitational observables are assumed to be asymptoticarkani ) that it is impossible to build a detector capable of probing the quantum theory of gravity. This might well be true but, as the next Section II argues, it begs questions which are directly associated with foundations of quantum mechanics rather than quantum gravity. On the other hand, as Section III describes, credible proposals not just of idealized detectors but of technologically achievable ones have been made which directly acces observables which are both quantum and gravitational. Thus, the conudrums described here might be solved in the scientifically best way, via experimental tests.
II Relationalism and unitarity
II.1 Bronstein’s argument and the quantum formalism
The main point hidden in Bronstein’s argument is that the usual “mathematical apparatus of quantum mechanics”, describing states as vectors in Hilbert space and observables as operators of this space, essentially assumes that the detector just “projects out the operator corresponding to the observable”, and in the process “collapses the wavefunction”. Many believe these fundamental assumptions of quantum formalism to be “physical”, and this belief is tied to the belief that quantum states and wavefunctions are “objective physical objects” (see for instance wigner ). In such a case, Bronstein’s argument implies that, provided the equivalence principle is exactly valid, there is new physics beyond quantum mechanics when gravitational effects are present (perhaps of the form penrose ; diosi ; curce ), whose mathematical formalization is wholly new, given the “tight” axiomatic structure of quantum theory witalg ; witalg2 ; vicary .
In fact, putting together recent results on Unruh detectors in a superposition of trajectories zych ; zych2 ; apadula (“quantum reference frames”) with the recent surprising argument wald ; wald2 that Rindler horizons collapse quantum superpositions, deviations from quantum mechanics if an ontological interpretation of its formalism is assumed become inevitable even for locally weak gravitational fields, for it would mean superimposed trajectories mean a superposition of timescales at which wavefunction collapse occurs, a paradox if such a collapse is a physical event.
However, others (generally ascribing to QBism, relationalism, the relative state and other ”informational” formulations) interp think that states and wavefunctions are nothing more than convenient mathematical objects to parametrize knowledge accessible to us. After all, non-commuting operators are inevitable if one assumes that some measurements are inherently incompatible with others; Furthermore, if our theory is to accomodate that
@@ II.1 Bronstein’s argument and the quantum formalism
Hermitian operators and Hilbert spaces then become the mathematical tool with the required characteristics.
The fundamental principle of this philosophy, as enunciated in rovelli , is that what we need for a description of the observable physical world is a physical description of physical systems relative to other systems, where relative means some kind of interaction. The axioms of quantum mechanics clearly fulfills these characterists, but more general axioms could also qualify, as for example open quantum systems described by non-unitary evolutions and with stochastic uncertainities as well as quantum ones. In this context, then, ”observables” are properties of the system relevant for the interaction with other systems. This is again a more general definition than that of usual quantum mechanics, with the latter arising in certain limits (the terms in Eq. 2 deriving from backreaction are negligible).
In this philosophy, the issues Bronstein raised are ”technical” rather than fundamental ones. One would just have to find more general mathematics objects, with better-suited characteristics in the regime where detector backreaction is significant.
II.2 Bronstein’s argument and relationalism
@@
Let us think what happens when the concepts above refer to coordinates, and one expects all quantum observables (i.e. correlators to all orders [^2] By definition, in a probabilistic theory all observables can be reduced to correlators of a certain order, $C(n)\sim\left\langle\phi(x_{1})...\phi(x_{n})\right\rangle$. Topological information can be encoded in $\lim_{n}C(n)$ and in $\lim_{x_{i}-x_{j}\rightarrow\infty}C(n)$) to be generally covariant w.r.t. coordinate changes frames . It will mean that
@@
It is now worth to compare the above generic definition of relationalism to the discussion, for example in witalg . The definition of rovelli given at the end of Section II.1 is entirely based on the axioms of probability combined with the limits of information extraction from a system inherent in quantum mechanics. From this, a set of relations are derived (Eq. 3 and Eq. 4) which look exactly like those required for a type I Von Neumann algebra. Eq. 5 relaxes this assumption by also considering the accessibility of information to the detector. In witalg the argument is made that generically one needs at least a type II algebra to describe observables in general, since a properly normalized density operator is needed to have well-defined averages and, eventually, compute the entropy. However a type I algebra (admitting pure states) is only needed for an observer which has full access to a complete time slice, not something that in a locally Lorentz invariant theory happens always. This means that the interpretation discussion in rovelli , when applied to infinite dimensional quantum fields, is fully consistent with the program in witalg to use the algebraic properties of quantum fields witalg2 to construct an algebra of background-independent observations accessible to an observer defined by a worldline. In the next Section II.3 we examine the physics consequences of this.
II.3 The equivalence principle instead of unitarity?
At first sight, the formulae in the previous section look like a trivial generalizations of the standard quantum formulae, but it is far from the case. In rovelli , the formalism is build around discreteness (ingrained in $N$ and the finite dimension of $U_{a,b}$). But coordinate transformations are continuus, necessitating a limit of $N\rightarrow\infty$ while $F$ become a function of the position operator $\hat{x}$ in quantum mechanics (one can define this via a lattice in configuration space). In quantum field theory, $F$ comes back to be a function of the label $x$ but $U$ becomes an operator valued function of the field operator $\hat{\phi}(x)$. As we know, the analytical properties of limits such as $N\rightarrow\infty$ can be different from any finite $N$ properties.
@@ II.3 The equivalence principle instead of unitarity?
associated with relativistic non-inertial frames. If the equivalence principle is exact, no contact with the regin beyond the horizon is possible. Hence,$F_{ij}$ is generally non-unitary and $U^{f},U^{f^{\prime}}$ are in different unitarity classes. Degrees of freedom beyond the horizon must be traced out by $F_{ij}$ for all questions, a la laflamme . Thus, transitivity as well as Eq. 4 fails and $F^{+}F\neq\delta$. Note that a time evolution constructed in a metric with a causal horizon should not be a unitary operator, but at most a mixing operator transforming covariantly under diffeomorphism shifts that exchange causal foliations. Note that this mixing operator can only be Lindbladian lind to leading order, its full non-Markovian structure would be determined by diffeomorphism constraints. As shown for instance in diosilind , a Lindbladian operator is not Lorentz invariant, while hu shows it will not have the right zero-temperature (Minkowski) limit.
@@
Thus, as described in witalg the observables accessible to a detector can not be described by a “wave-function” vector in Hilbert space if Eq. 5 holds. The tension between the definition of local observables, unitarity and diffeomorphism invariance has already been noted in works such as hao1 ; hao2 . If however these observables form a type II algebra, as witalg argues for, they can be described by a “wave-function” as a parametrization of a non-pure density matrix, $\hat{\rho}_{ij}=\left|\Psi\right>_{i}\prescript{}{j}{\left<\Psi\right|}$ (but since there is no Hermitian Hamiltonian there are also no pure states commuting with it) or equivalently a partition function nishioka (where general convariance is imposed both on the action and the domain of functional integration, creating a ”holographic” constraint holo ). The question of what substitutes unitarity as a constraint on operators in the presence of general covariance is of course very much open. In megrav it is speculated that for general covariance unitarity $\left(\mathrm{Re}U+\mathrm{Im}U=\left(\mathrm{Re}U-\mathrm{Im}U\right)^{-1}\right)$ is relaxed to a general Kramers-Konig type tong fluctuation dissipation relation $\left(\begin{array}[]{cc}\mathrm{Re}\\ \mathrm{Im}\end{array}U=\begin{array}[]{c}F_{1}(\mathrm{Im}U)\\ F_{2}(\mathrm{Re}U)\end{array}\right)$ for some $F_{1,2}(...)$ determined locally by dynamics, where the ”Real” fluctuation term dominates for weak fields (and represents the equivalence between coordinate metric superpositons and detector superpositions) while the ”Imaginary” dissipation dominates for stronger fields and represents decoherence from horizons, a la wald . In the operator picture, a dynamically determined fluctuation dissipation relation reflects the well-known fact that the preparation of mixed quantum states is not unique preskill . In this respect the local “shaking” of the detector corresponds to incoherent sums of position eigenstates ($\rho\rightarrow\hat{\lambda_{i}}|x_{i}(t)><x_{i}(t)|$ where $\hat{\lambda}$ are classical probabilities of the amount of shaking and $|x_{i}(t)><x_{i}(t)|$ the detector’s position in its co-moving frame ) while horizon decoherence represents tracing out of DoFs in the configuration space basis. In principle, one can see how the two effects could compensate; In other words, if every state is inherently mixed, one could enforce diffeomorphism invariance by demanding that every diffeomorphism represents a state prepatation representing the same density matrix for the causally accessible region by the observer.
This means that Dirac type quantization (behind relations such as Eq. 1) should not anymore apply: Eq. 1 is a consequence of Dirac quantization relations woit , where canonical pairs of observables have a commutator $i\hbar\delta$ ( The Baker–Campbell–Hausdorff formula implies such a commutator for the Stone-Von Neumann theorem to apply). Instead, we expect that uncertainity relations to have a component based on commutators (which is related to the uncertainity of another operator), on the detector backreaction (related to the detector’s properties) and on the observable’s accessible horizon (which is a ”minimum” bound not related to any other operator, similar to a thermal uncertainity, and in fact interpretable as one via a horizon temperature).
@@
II.4 A generally covariant ”quantum” theory: the technical challenge
@@
Thus, the problem of formulating a theory which is both quantum and obeys the equivalence principle exactly has been reformulated as a problem of finding a generally covariant representation for $F,U$ in Eq. 5. This is of course a very difficult problem ($F,U$ are infinite dimensional, and $F$ represent coordinate diffeomorphisms). So far it is fair to say no such theory is conclusively proven to exist. The only candidates of quantum gravity explicitly built around the fully quantum equivalence principle, Loop Quantum Gravity and causal dynamical triangulation lqg ; cdt identify spacetime coordinates (trough area operators,locally quantized via an SU(2) symmetry but transforming covariantly under diffeomorophisms) with observables (so there is no separation between $U$ and $F$) and impose general covariance on the state, something that is different from the relational discussion in this work, centered around observables rather than states [^3] A direct consequence of diffeomorphism invariance of the quantum state in LQG is the reduction of Unitary time evolution to $\hat{H}|\psi>=0$. We note that the difficoulty of defining time dynamically has given rise to an associated interpretation of quantum measurement montevideo , based on assuming any time measurement as involving decoherence wootters ; rovelli . The role of general covariance at the quantum level in this interpretation is not clear. In realistic physical experiments we do not measure coordinates directly, but infer them via interactions (“rulers” and “clocks”), hence the distinction between $U$ and $F$. General covariance means not the covariance of a “state” (As can be seen, for instance in kovner in Non-Abelian Gauge theories wave functionals are explicitly not Gauge-invariant) but of observable correlations (matrices $\left|U_{ab}\right|^{2}$, whose basis are field correlators, under $F-$ type transformations), and general covariance for field correlators living on a quantum spacetime is not as yet proven for LQG or CDT lqg ; cdt . What this means for our discussion is that while it is likely that the Algebra inherent in lqg ; cdt is both of type I and generally covariant, it is not clear what algebra one would get, and if such an algebra would be generally covariant, if one combines the spacetimes of lqg ; cdt with a field and an observer making observations in such a field.
It is also worth to remark that, to tree level in certain limits quantum field theory already fulfills the requirement of general covariance matsas so the hope that this can be brought to higher orders in an EFT expansion, as suggested by megrav ; holo is not unfounded. This approach entails reconstructing the metric from QFT observables and seeing what sort of QFT action would lead to a general covariance in terms of this of metric. Such investigations are speculative but ongoing (eg padma ; kempf1 ; kempf2 ; tales ; nist ).
@@
Most importantly, the problem defining this Section II.4 can be thought of as ”technical” rather than ”conceptual”. In fact procedures for achieving this have already been proposed fewster . The price for this problem reduction is itself however both technical and conceptual in nature. The technical price is the abandonment of the very convenient Hilbert space and Unitary/Hermitian operator technology. The conceptual price is a definite abandonment of notions that wavefunctions and states are physical objects. General covariance makes it inevitable that the only mathematical objects readily associated with a physical meaning are “sets of questions that can be asked of a system”, i.e. density operators in some observable’s basis. For anyone who does not associate an ontic truth to quantum states, this is a small price to pay, or equivalently a feature rather than a bug.
III An experimental perspective
III.1 Testing the quantum nature of gravity experimentally
@@
III.2 The Bose-Marletto-Vedral experiment and quantum reference frames
Looking at the experiment described in the previous section from the point of view of quantum reference frames frames indeed confirms that the analysis in witness ; witness2 ; gyro ; gyro2 ; pend1 is done in the laboratory frame. Using the spin set-up in witness , Eq. 5’s $U$ will be in the basis of all possible $\uparrow_{1}\uparrow_{2}(x_{in})\right|\left|\uparrow_{1}\downarrow_{2}(x_{out})\right>$ where $\uparrow_{1,2}$ transforms as a spinor under the Lorentz group. Provided the nanoparticles being and end in the same place (something not explicitly required in witness ; witness2 but can be easily accomplished, as Fig. 1 shows) the phase change is the difference between two angles propagated through two proper time intervals. Thus in the classical limit binney given a unit vector $A^{\mu}$ the phase difference $\sim(\Delta A_{\mu})A^{\mu}$ where $\Delta A^{\mu}=R^{\mu}_{\alpha\beta\gamma}\Delta S^{\alpha\beta}A^{\gamma}$ ($A^{\mu}$ is the vector projection of the difference between two spinors, a vector), where $S^{\alpha\beta}$ is the area element parallel to the interferoemeter and $R^{\mu}_{\alpha\beta\gamma}$ the Riemann tensor of the gravitational field (sum of earth’s, apparatus and nanoparticles). If tidal forces were to be ignored (they can for the earth, but not the rest) this would of course be a 4-vector. We are looking, therefore, for the observable to transform covariantly as a 4-vector if general covariance works in the quantum regime.
@@ III.2 The Bose-Marletto-Vedral experiment and quantum reference frames
IV Conclusion
The only conclusion we can make with some confidence given the conudrums outlined in section II, and the proposals in section III is that the study of quantum gravity might be in it’s ”golden” phase: Deep conceptual issues in both gravity (the role of the equivalence principle) and quantum mechanics (fundamental limits to the “classicality” of the detector and the “quantumness of the system) might be directly tested by experiments of the type witness ; witness2 ; gyro ; gyro2 ; pend1 . If the answer estimated by witness will be seen, one can interpret this as a confirmation of the quantization of gravity but also an experimentally verified breakdown of general covariance of quantum observables (or rather of the vacuity of the concept of general covariance in the quantum regime), as advocated in dono . If the result of witness ; witness2 ; gyro ; gyro2 ; pend1 can not be described by canonical quantum mechanics but rather with some kind of mixed theory a la oppenheim ; megrav ; holo and at the same time stringent tests of the equivalence principle with both astrophysical observations nordveldt and tabletop measurements atoms yield no violation, one would have to think deeper at how to implement the equivalence principle into a “relational generally covariant Algebra of observables” woit ; witalg , which in general will not be based around unitary matrices. The advent of mesoscopic experimental study of quantum effects on gravity, therefore, promises to be important and fruitful.